HC Verma chapter Centre of mass, linear momentum, collision Pg163 Question 51Q two blocks of Masses m1 and m2 are connected by a spring of spring constant k. Suppose each of the block is pulled by constant force F. Find the maximum elongation That the spring will suffer.

Grade:11

1 Answers

Vikas TU
7 years ago
From the work energy Theorem,
F*x + F*x = 0.5kx^2
2Fx = kx^2/2
2F = kx/2
x = 4F/k .
would be the maximum elongation.